OpenStudy (anonymous):

Solve the following system: x - y + z = 0 3x - 2y + 5z = 8 2x + y - 3z = -9

OpenStudy (anonymous):

id suggest you pick 1 variable on the first line, maybe x, and rewrite it. x = y - z. now plug this is for x in the 2nd and 3rd row. this will eliminate x completely. you can then do the same concept as you did for x for either y or z in the 2nd row, to be done in the 3rd

OpenStudy (anonymous):

Another possibility is to multiply the third equation by 2 and add it to the second to get an equation in x and z. Then multiply the first equation by 2 and subtract it from the second to get another equation in x and z. U now have simultaneous equations in x and z which you can solve normally. Sub back to get y.

OpenStudy (anonymous):

x - y + z = 0 ----------(1) 3x - 2y + 5z = 8 ----------(2) 2x + y - 3z = -9 ----------(3) first use (1) and (2) to eliminate x x - y + z = 0 ----------(1) multiply by -3 3x - 2y + 5z = 8 ----------(2) -3x + 3y - 3z = 0 3x - 2y + 5z = 8 add these two to get y + 2z = 8 ---------------(4) now use (2) and (3) to eliminate x 3x - 2y + 5z = 8 ----------(2) multiply by 2 2x + y - 3z = -9 ----------(3) multiply by -3 6x - 4y + 10z = 16 -6x - 3y + 9z = 27 add these two to get - 7y + 19z = 43 --------------(5)

OpenStudy (anonymous):

more following...

OpenStudy (anonymous):

now take (4) and (5) to eliminate y y + 2z = 8 --------------(4) multiply by 7 -7y + 19z = 43 --------------(5) 7y + 14z = 56 -7y + 19z = 43 add these two to get 33z = 99 z = 99/33 = 3 so z = 3 <<<<< substitute z = 3 in y + 2z = 8 --------------(4) to get y + 2(3) = 8 y = 8 - 6 y = 2 so y = 2 <<<<< substitute y = 2 and z = 3 in x - y + z = 0 ----------(1) x -(2) + (3) = 0 x - 2 + 3 = 0 x + 1 = 0 x = -1 so x = -1 <<<<< so solution is (x, y, z) = (-1, 2, 3)

OpenStudy (anonymous):

i hope the method is clear .....☺